the sum of the digits of a two-digit number is 14. if the digits are reversed, the number increased by 18. find the number.
Answer by mananth(16946)

the sum of the digits of a two-digit number is 14. if the digits are reversed, the number increased by 18. find the number.
..
let digits be x & y x in ten's place
x+y=14.............1
...
10y+x= 10x+y+18
9y-9x=18
/9
y-x=2.............2
add (1)&(2)
2y=16
/2
y=8
so x= 6
so the number is 68
...
m.ananth@hotmail.ca